# NGO Activism:

### v6

@@ Abstract
We examine how NGOs time campaigns to influence corporate decisions at Annual General Meetings (AGMs). Using data on 2,500 U.S. campaigns, we find that NGOs frequently launch campaigns on the targeted firm’s AGM date, which boosts media exposure and triggers future shareholder proposals but does not sway current AGM votes or affect sales. As NGOs gain experience and reputational capital, they shift to earlier pre-AGM campaigns that increase shareholder support at the AGM and lead to improvements in corporate behavior. These patterns reveal a strategic trade-off between maximizing exposure and achieving direct influence.
@@ 1 Introduction
Examining campaign timing reveals that less experienced NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020). These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents. For instance, as the NGO Public Eye put it in its 2025 campaign timed on the AGM date of Inditex: “As shareholders, you have the power to put Inditex on a sustainable course” (Clean Clothes Campaign, 2025). Indeed, we find little evidence that other stakeholders are moved by these campaigns. For example, we find no evidence that sales are differently affected based on campaign timing relative to the AGM date, indicating that consumers may not be the main audience of NGOs.
@@
While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves. Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively inexperienced. These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times. This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
@@
As NGOs accumulate visibility through AGM-date campaigns, their strategic calculus changes. Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes. By contrast, campaigns launched in the months leading up to the AGM substantially raise the likelihood that related shareholder proposals succeed. Intervening one to three months before the meeting boosts average shareholder support by 8%, while even earlier campaigns often prompt firms to settle before proposals reach a vote, leading to measurable improvements in environmental and social performance.
@@
Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns forward in time as they gain experience and reputational capital. First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions. This pattern is especially pronounced among NGOs with stronger reputations, proxied by higher search visibility, which campaign several months ahead of the AGM rather than on the day itself.
Repeated targeting of AGMs also plays a distinct role. By returning to these focal moments, NGOs build credibility and reputational standing, which in turn increases the effectiveness of later campaigns. Yet such returns are not indefinite: once credibility is established, additional AGM-day campaigns yield diminishing benefits. Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—shows how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for firms’ social and environmental impact.
In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982). As NGOs accumulate reputational capital (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their strategy shifts: campaigns are launched months before the AGM, when they can still sway shareholder support and discipline managers.
@@
By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022). More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
@@ 4 Timing Campaigns for Visibility
Under the hypothesis that AGM campaigns affect the visibility of experienced and inexperienced NGOs differently, we estimate (1) separately for an NGO’s first few campaigns and for later ones. Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and its subsequent campaigns (red triangles). The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months. For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established. These patterns suggest that inexperienced NGOs rely more heavily on AGM timing to build name recognition, whereas established NGOs gain less from repeating the same tactic.
@@
We show that these results are robust to addressing three main concerns. First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months. This misalignment can distort estimates when treatment effects are heterogeneous across groups or over time. Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other. To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year. Both approaches yield similar results, with much larger gains for “unexperienced” NGOs.
@@
Further evidence on selection based on NGO experience comes from Table A1. Panel (c) compares firms targeted by NGOs with high versus low campaign experience and finds no significant differences across the same variables studied in Panel (b) of Figure 2. Panel (d) corroborates this by regressing firm characteristics on NGO experience (i.e., number of past campaigns), showing no correlation once fixed effects are included. The only systematic trend is that NGOs increasingly focus on environmental rather than social issues in later years, reflecting their rising importance.
@@
Gains in visibility are greater for unexperienced NGOs campaigning on the target’s AGM day than for experienced NGOs.
@@ 4.1 Stakeholders’ Reactions
We next ask how campaign timing matters for stakeholders and firms’ social and environmental outcomes. We focus on shareholders and consumers, with supporting exhibits in the Online Appendix.
@@ 4.1.2 Consumers
Appendix Figure B4 shows the estimated $\hat{\beta}_{\tau}$} coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns. [^5] Unreported results show also no difference between experienced and unexperienced NGOs.
@@ 4.2 Does AGM-Targeting Create Positive Change?
If public shaming effectively influenced firm behavior, the AGM month should deliver the largest improvements in E&S scores, given that it triggers heightened negative news coverage of the target firm (Appendix Figure B2). However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year. By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
Although ESG scores are an imperfect measure of a firm’s social impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), this finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore, the results in this section suggest that:
@@ 5 Timing Campaigns to Influence AGM Outcomes
Sections 4 and 5 outlined a tradeoff: campaigns launched close to an AGM maximize exposure, while those launched earlier maximize influence. In this section, we examine how NGOs navigate this tradeoff as their visibility increases. We first study whether greater experience or reputation leads NGOs to shift campaign timing from the AGM date to earlier months. We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
6.1 NGO Experience and Campaign Timing
We first examine whether NGOs, as they gain experience, shift from targeting AGMs to launching earlier campaigns. Specifically, we estimate
@@ 6.1 NGO Experience and Campaign Timing
These patterns are not driven by alternative explanations. First, they are unlikely to reflect NGOs learning that campaigning on the AGM date does not affect voting, since most shareholders have already voted by proxy. Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7). Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO experience (Appendix Figures B8 and B9).
6.2 Reputation and Campaign Timing
@@
Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation or experience grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
@@
NGOs initially target AGM dates to gain visibility. As their experience or reputation increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
@@
(a) More experienced NGOs campaign earlier
@@ 6.3 NGO Reputation and AGM Influence
We now ask whether NGO experience and reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{Experience}_{g,y-1}$ (7)
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{Experience}_{g,y-1}$
@@
where $\text{Success}_{i,n,y}$” indicates whether a proposal is withdrawn or passes a vote, as in equation (4). The key variable $\text{Campaign on Same Cause}_{i,n,y}$” equals one if NGO $g$ campaigned against firm $n$ on cause $i$ in the ten months preceding the AGM. $\text{Experience}_{g,y-1}$” counts NGO $g$’s past campaigns. We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
Table 1: NGO experience at targeting AGMs and success at supporting current shareholder proposals
@@
Experience -0.034
Campaign on Same Topic $\times$ Experience 0.078∗∗∗
Experience square 0.008∗∗
Campaign on Same Topic $\times$ Experience square -0.015∗∗
Non-AGM Experience 0.0008
Campaign on Same Topic $\times$ Non-AGM Experience -0.011
AGM Experience 0.005 -0.034
@@
Campaign on Same Topic $\times$ AGM Experience 0.048 0.091∗∗∗
@@
AGM Experience square 0.010∗∗
Campaign on Same Topic $\times$ AGM Experience square -0.013∗∗
@@
  • Note: This table reports the estimated coefficients from (7) in Column 1. In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (Non-AGM Experience) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), and their interaction with Campaign on Same Topic. In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (AGM Experience), its square, and their interaction with Campaign on Same Topic. In Column 4, we replace Experience from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign. All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects. Clustered (firm) standard errors are reported in parentheses.
@@
Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$). What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has accumulated experience ($\beta_{3}>0$). A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average.
@@
Column (2) distinguishes between campaigns held on AGM dates and those held at other times. Only the former contribute to proposal success. In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
@@
Column (3) tests for diminishing returns by including squared terms. The results confirm a concave pattern: additional AGM campaigns continue to improve proposal success, but at a decreasing rate. Importantly, the combined effect remains non-negative even for the most active NGOs. [^8] The marginal effect of AGM Experience and its square never turns negative, so our proxy for reputation does not imply adverse returns.
@@
We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends. Specifically, Column (4) replaces experience by Google Search Popularityg,y, which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.999The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects. The interaction between this measure and campaigning on the same cause is positive, with a concave shape. Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
These results also resonate with the media evidence in Section 4: AGM targeting increases negative media coverage only for relatively inexperienced NGOs. More seasoned NGOs gain little additional visibility from campaigning exactly on AGM dates and may be more effective when campaigning slightly earlier. These results support our view that:
@@
Targeting AGMs builds NGO reputation and increases their ability to raise the supported of related shareholder proposals, but with diminishing returns.

### v7

@@ Abstract
We examine how NGOs time campaigns to influence corporate decisions at Annual General Meetings (AGMs). Using data on 2,500 U.S. campaigns, we find that NGOs often campaign on the AGM date, which increases media exposure and future shareholder proposals but does not affect current AGM votes or consumers. As NGOs gain reputational capital, they shift to earlier pre-AGM campaigns to support related shareholder proposals and directly improve corporate behavior. The evidence underscores a strategic trade-off between visibility and influence, showing how NGO tactics evolve with NGOs’ credibility and the changing expectations of socially conscious stakeholders about NGO activities.
@@ 1 Introduction
Examining campaign timing reveals that younger NGOs concentrate campaigns on the AGM date, where visibility is maximized but most shareholder votes are already casted (Gent et al., 2015; Bebchuk et al., 2020). These campaigns cannot change immediate voting outcomes but are found to mobilize other stakeholders, such as future shareholder proponents. For instance, as the NGO Public Eye put it in its 2025 campaign timed on the AGM date of Inditex: “As shareholders, you have the power to put Inditex on a sustainable course” (Clean Clothes Campaign, 2025). Indeed, we find little evidence that other stakeholders are moved by these campaigns. For example, we find no evidence that sales are differently affected based on campaign timing relative to the AGM date, indicating that consumers may not be the main audience of NGOs.
@@
While they do not sway votes or sales, AGM-timed campaigns significantly increase the visibility of NGOs themselves. Using Google search data, we find that such campaigns generate sharp increases in search intensity for the campaigning NGO, particularly when it is relatively young. These visibility gains persist for several months and are substantially larger than those observed for campaigns launched at other times. This suggests that AGM-day actions, while limited in shifting immediate corporate outcomes, are instrumental in boosting public recognition of NGOs.
@@
As NGOs accumulate visibility through AGM-date campaigns, their strategy changes. Because most shareholder votes are cast well before the meeting, campaigns on the day itself rarely sway immediate outcomes. By contrast, campaigns in the months preceding the AGM raise the likelihood of proposal success. Campaigns one to three months before the meeting increase support by 8%, while earlier ones often induce settlements that improve firms’ environmental and social performance.
@@
Consistent with this shift from exposure to influence, we find that NGOs progressively move their campaigns earlier in time as they gain AGM exposure. First-time campaigns are concentrated on the AGM date, where visibility is highest, while subsequent efforts are launched progressively earlier, when they can still shape shareholder decisions. This pattern is especially pronounced among NGOs with high search intensity, which may not drive substantial visibility benefits from AGM-day campaigns.
Repeated targeting of AGMs also plays a distinct role. By returning to these focal moments, NGOs build a “reputation for toughness”—demonstrating their ability to amplify negative media coverage when firms are most exposed—which can enhance the effectiveness of subsequent campaigns. Yet, once reputation is established, further AGM-day campaigns yield diminishing benefits, generating smaller increases in NGO search intensity, lower success rates for related shareholder proposals, and less negative media coverage on the target firm. This pattern reflects NGOs’ changing incentives because of evolving stakeholder expectations. As NGOs become recognized and credible actors (Bar-Isaac et al., 2008; Baur and Palazzo, 2011; Mitchell and Stroup, 2017), their actions are interpreted differently: what initially signals resolve later signals routine. Established NGOs are therefore expected to engage in more constructive tactics, such as supporting shareholder proposals, rather than relying on visible confrontation.
Taken together, this lifecycle of tactics—initially focused on visibility, later on direct influence—illustrates how NGOs adaptively balance exposure, reputation, and pressure on firms, with implications for their social and environmental impact. In this sense, AGM-day campaigns act as costly signals of commitment that help NGOs build credibility (Kreps and Wilson, 1982; Milgrom and Roberts, 1982), but the meaning and effectiveness of those signals evolve as stakeholders update their expectations about each NGO’s toughness and legitimacy.
@@
By documenting this strategic adaptation, we provide new conditions under which firms internalize their externalities (Fioretti, 2022; Allcott et al., 2022). More broadly, our work offers empirical support for theories asserting that stakeholders contribute resources only when they expect to capture part of the value created, underscoring the interplay between stakeholder involvement, reputation, and corporate change (Porter and Kramer, 2011; Kaplan, 2019; McGahan and Pongeluppe, 2023).
@@ 4 Timing Campaigns for Visibility
Under the hypothesis that AGM campaigns affect NGO visibility differently depending on their reputation, we estimate (1) separately for an NGO’s first few campaigns, when its recognition is still limited, and for later ones. Panel (a) of Figure 2 reports the estimated coefficients of $\beta^{AGM}_{\tau}$, splitting the sample into an NGO’s first three campaigns (blue dots) and subsequent campaigns (red triangles). The results indicate that AGM campaigns generate a sharp and statistically significant increase in NGO visibility, raising Google Trends scores by almost 50% in the following three months. For later campaigns, the effect is weaker—about 40% in the first month and around 10% in the next two months—and not statistically significant in any period, consistent with diminishing visibility returns as NGOs become more established. These results are estimated within NGOs: early in their lifecycle, NGOs benefit more from AGM timing to build visibility, but the returns to this tactic decline as their public recognition grows.
@@
We show that these results are robust to addressing three main concerns. First, the stacked event-study framework implicitly treats campaign timing as simultaneous across NGOs, whereas in reality campaigns are staggered across organizations and calendar months. This misalignment can distort estimates when treatment effects are heterogeneous across groups or over time. Second, some NGOs launch multiple campaigns within the same year, which may confound the analysis if two campaigns fall within five months of each other. To address both issues, Appendix C replicates the analysis using the staggered difference-in-differences estimator of Sun and Abraham (2021), which explicitly accounts for variation in treatment timing, as well as a distributed-lag specification that flexibly controls for past campaigns within a year. Both approaches yield similar results, with substantially larger gains for early campaigns.
@@
Further evidence on potential selection based on the number of past NGO campaigns is presented in Table A1. Panel (c) compares firms targeted by NGOs that had previously campaigned on a target’s AGM date with those that had not, finding no significant differences across the same variables analyzed in Panel (b) of Figure 2. This suggests that NGOs do not appear to select targets based on their past campaigning strategies. Panel (d) corroborates this result by regressing firm characteristics on the number of past NGO campaigns and showing no systematic relationship once fixed effects are included. The only consistent pattern is that NGOs increasingly focus on environmental rather than social issues over time, reflecting their growing salience.
@@
Visibility gains from AGM‐day campaigns are greatest during an NGO’s initial campaigns.
@@ 4.1 Stakeholders’ Reactions
We next examine how campaign timing affects stakeholders and firms’ ESG outcomes.
@@ 4.1.2 Consumers
Appendix Figure B4 shows the estimated $\hat{\beta}_{\tau}$} coefficients separately for NGO campaigns (i) planned on an AGM date (Panel a) and (ii) not planned on an AGM date (Panel b). We find no consumer response in either case, indicating that shareholders may be the primary audience for NGO campaigns. [^5] Unreported results show also no difference between early and late campaigns.
@@ 4.2 Does AGM-Targeting Create Positive Change?
If public shaming effectively influenced firm behavior, the AGM month should deliver the largest improvements in E&S scores, given that it triggers heightened negative news coverage of the target firm (Appendix Figure B2). However, Appendix Figure B5 shows that only campaigns launched 3 to 4 months before the AGM lead to subsequent increases in firms’ E&S scores—rising by around 3% over the following year. [^6] Acknowledging the limits of ESG scores as measures of social and environmental impact (e.g., Kotsantonis and Serafeim, 2019; Allcott et al., 2022), we use MSCI to mitigate signal risk: its ratings are more time-consistent and widely used by sustainable investors (Berg et al., 2021, 2024). By contrast, campaigns launched at other times, including during the AGM month, have no economically or statistically significant effect.
This finding suggests that NGOs’ preference for targeting AGM dates may be driven by factors beyond their direct influence on firms. Therefore, the results in this section suggest that:
@@ 5 Timing Campaigns to Influence AGM Outcomes
Sections 4 and 5 outlined a tradeoff: campaigns launched close to an AGM maximize exposure, while those launched earlier maximize influence. In this section, we examine how NGOs navigate this tradeoff as their visibility increases. We first study whether greater past exposure leads NGOs to shift campaign timing from the AGM date to earlier months. We then ask whether such reputation translates into greater effectiveness at influencing shareholder proposals.
6.1 The Timing of Past and Current Campaigns
We first examine whether NGOs shift from targeting AGMs to launching earlier campaigns as they campaign more. Specifically, we estimate
@@ 6.1 The Timing of Past and Current Campaigns
These patterns are not driven by alternative explanations. First, they are unlikely to reflect NGOs learning that campaigning on the AGM date does not affect voting, since most shareholders have already voted by proxy. Indeed, the same downward trend in campaign timing holds when focusing on NGOs whose first campaign occurred after the AGM (Appendix Figure B7). Second, they are not driven by selection of more “influenceable” targets: neither target size (market capitalization) nor ESG scores covary with NGO past AGM exposure (Appendix Figures B8 and B9).
6.2 Visibility and Campaign Timing
@@
Taken together, these results indicate a systematic shift: NGOs start by targeting AGMs to maximize exposure, and as their reputation grows, they move campaigns forward in time to increase their chances of influencing firm decisions.
@@
NGOs initially target AGM dates to gain visibility. As their exposure increases, they advance campaigns by several months to influence the outcome of shareholder proposals.
@@
(a) Later campaigns pre-date the target’s AGM date
@@ 6.3 NGO Reputation and AGM Influence
We now ask whether NGO reputation, as measured in the previous sections, translates into greater influence over shareholder proposals. To test this, we estimate
$\text{Success}_{i,n,y}$ $=\beta_{1}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}+\beta_{2}\,\text{\# Past Campaigns}_{g,y-1}$ (7)
$+\beta_{3}\,\text{Campaign on Same Cause}_{c(i,n,g,y,m)}\cdot\text{\# Past Campaigns}_{g,y-1}$
@@
where $\text{Success}_{i,n,y}$” indicates whether a proposal is withdrawn or passes a vote, as in equation (4). The key variable $\text{Campaign on Same Cause}_{i,n,y}$” equals one if NGO $g$ campaigned against firm $n$ on cause $i$ in the ten months preceding the AGM. “# $\text{\# Past Campaigns}_{g,y-1}$” counts NGO $g$’s past campaigns. We absorb firm-by-year, NGO-by-year, cause-by-year, and month variation through fixed effects, and cluster standard errors at the firm level.
Table 1: NGO’s strategy and support for related shareholder proposals
@@
# Past Campaigns -0.034
Campaign on Same Topic $\times$ # Past Campaigns 0.078∗∗∗
# Past Campaigns squared 0.008∗∗
Campaign on Same Topic $\times$ # Past Campaigns squared -0.015∗∗
# Past Non-AGM Campaigns 0.0008
Campaign on Same Topic $\times$ # Past Non-AGM Campaigns -0.011
# Past AGM Campaigns 0.005 -0.034
@@
Campaign on Same Topic $\times$ # Past AGM Campaigns 0.048 0.091∗∗∗
@@
# Past AGM Campaigns squared 0.010∗∗
Campaign on Same Topic $\times$ # Past AGM Campaigns squared -0.013∗∗
@@
  • Note: This table reports the estimated coefficients from (7) in Column 1. In Column 2, we also consider the number of campaigns the NGO has launched in the past which were not on the firm’s AGM day (# Past Non-AGM Campaigns) and the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), and their interaction with Campaign on Same Topic. In Column 3, we focus on the number of campaigns launched by the NGO on the targeted firm’s AGM date in the past (# Past AGM Campaigns), its square, and their interaction with Campaign on Same Topic. In Column 4, we replace # Past Campaigns from model 7 by NGO $g$’s average Google Trends Score relative to GreenPeace the year before the campaign. All regressions also include firm-by-year, NGO-by-year, topic-by-month-by-year fixed effects. Clustered (firm) standard errors are reported in parentheses.
@@
Column (1) of Table 1 shows that campaigning on the same cause as a shareholder proposal has little effect on its own ($\beta_{1}\approx 0$). What matters is the interaction: proposals are more likely to succeed when the campaigning NGO has already campaigned ($\beta_{3}>0$). A one-standard-deviation increase in past campaigns on the same cause raises proposal success by about 8% on average. The regression also includes the squared term of “# $\text{\# Past Campaigns}_{g,y-1}$,” showing decreasing marginal returns of campaigning.
@@
We unpack the mechanism behind these coefficient estimates by distinguishing between prior AGM campaigns and non-AGM campaigns in Column (2). We find that only the former contribute to proposal success. In fact, one-standard-deviation more AGM campaigns increases the probability of success by roughly 4.5 percentage points (about 15%), while non-AGM campaigns have no measurable effect.
@@
Column (3) tests for diminishing returns by including squared terms. The results confirm a concave pattern: additional AGM campaigns continue to improve proposal success, but at a decreasing rate. Importantly, the combined effect remains non-negative even for the most active NGOs. [^9] The marginal effect of AGM campaigning never turns negative in our sample since no firm launches more than five AGM-date campaigns, so our proxy for reputation does not imply adverse returns.
@@
We corroborate these findings with an alternative proxy: NGO visibility measured by Google Trends. Specifically, Column (4) replaces “# $\text{\# Past Campaigns}_{g,y-1}$” by “Google Search Popularityg,y-1,” which we define as the average search score of NGO $g$ relative to Greenpeace in year y1y-1.101010The table drops the direct effect of this variable on the dependent variable, as it is absorbed by NGO-by-Year fixed effects. The interaction between this measure and campaigning on the same cause is positive, with a concave shape. Thus, both proxies—past AGM campaigns and Google visibility—suggest that reputation enhances the effectiveness of NGO campaigns, albeit with diminishing marginal returns.
These results support the following result:
@@
Targeting AGMs builds NGO reputation and increases their ability to raise the support of related shareholder proposals, but with diminishing returns.